An airplane travels from points A and B with a distance of 1500 km and a wind along its flight line. If it takes the airplane 2 hours from A to B with the tailwind and 2.5 from B to A with the headwind, what is the velocity?
Answer by checkley75(3666)

1500/2=750 KMH
1500/2.5=600 KMH
750-W=600+W
2W=750-600
2W=150
W=150/2
W=75 KMH FOR THE WIND.
750-75=675 FOR THE AIRPLANE SPEED.